OpenStudy (anonymous):

For the line with slope -2 and passing through (3,-2) write the equation of the line in: a) point slope form b) slope form c) standard form

OpenStudy (whpalmer4):

point-slope form: \[y - y_1 = m(x-x_1)\]where \(m\) is slope and \((x_1,y_1)\) is a point the line goes through slope form (probably means slope-intercept form): \[y = mx + b\] where \(b\) is the y-value of the y-intercept You can get this from the equation for point-slope form simply by solving for \(y\) standard form: \[Ax+By=C\]where \(A>0\), \(A, B, C\) are integers if possible, and relatively prime (in other words, you can't factor out anything)